OpenStudy (anonymous):

Mike shoots a large marble (Marble A, mass: 0.05 kg) at a smaller marble (Marble B, mass: 0.03 kg) that is sitting still. Marble A was initially moving at a velocity of 0.6 m/s, but after the collision it has a velocity of –0.2 m/s. What is the resulting velocity of marble B after the collision? Be sure to show your work for solving this problem along with the final answer.

OpenStudy (shamim):

Use m1*u1+m2*u2=m1*v1+m2*v2

OpenStudy (shamim):

Given m1=0.05kg m2=0.03kg

OpenStudy (shamim):

u1=0.6m/s u2=0

OpenStudy (shamim):

v1=-0.2m/s v2=?